# DAPS++: Rethinking Diffusion Inverse Problems with Decoupled Posterior Annealing

### v1

@@ 5 Conclusion
In this work, we revisit the underlying mechanism of diffusion-based inverse problem solvers and propose an expectation–maximization (E–M) style formulation that makes explicit the separation between the diffusion and data-consistency stages. Based on this perspective, we develop DAPS++, a fully decoupled framework that substantially improves sampling efficiency while preserving high reconstruction fidelity. The diffusion stage functions as a prior-driven initialization on a flat manifold, whereas the data-consistency stage performs likelihood-guided refinement independent of the diffusion process. This separation offers a clearer Bayesian interpretation of diffusion-based inference and provides an explanation for the empirical behavior of existing approaches. Across both linear and nonlinear imaging tasks, DAPS++ achieves strong performance using only a fraction of the function evaluations required by conventional methods, establishing an efficient and scalable foundation for future research in diffusion-based inverse problems.
[…]
@@
6 Theoretical Analysis
6.1 Lipschitz Analysis of Prior–Likelihood Interaction
@@
$\mathbf{r}_{t}:=\mathbf{y}-\mathcal{A}\,\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}).$ (11)
@@
$\kappa_{t}\;\geq\;\frac{\sigma_{\min}^{+}(\mathcal{A})}{\gamma^{2}C}\,\sigma_{t}\,\|\mathbf{r}_{t}\|.$ (12)
@@
$\|\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t};\sigma_{t})\|\;\lesssim\;\frac{C}{\sigma_{t}}.$ (13)
@@
$\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}\mid\mathbf{x}_{t};\sigma_{t})=\frac{1}{\gamma^{2}}\mathcal{A}^{\top}\mathbf{r}_{t}+O(\sigma_{t}).$ (14)
@@
$\|\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}\mid\mathbf{x}_{t};\sigma_{t})\|\;\geq\;\frac{\sigma_{\min}^{+}(\mathcal{A})}{\gamma^{2}}\|\mathbf{r}_{t}\|.$ (15)
(iii) Lower bound on gradient ratio. Combining (13) and (15),
$\kappa_{t}\;\gtrsim\;\frac{\sigma_{\min}^{+}(\mathcal{A})}{\gamma^{2}C}\,\sigma_{t}\,\|\mathbf{r}_{t}\|,$ (16)
establishing (12). $\square$
@@
This analysis rigorously justifies the observation that the prior gradient contributes negligibly in high-noise regimes. For standard EDM schedulers in DAPS with $\sigma_{\max}\approx 100$ and low observation noise ($\gamma\approx 0.01$), the gradient ratio reaches magnitudes of $\kappa_{t}\approx 10^{6}$, rendering the prior term mathematically insignificant during early sampling. As illustrated in Fig. 1, this dominance persists throughout the diffusion process, where the likelihood gradient continues to outweigh the prior contribution even at the minimal noise levels ($\sigma_{\min}\approx 0.1$). Consequently, even for ill-posed problems where $\mathcal{A}$ has a null space, the likelihood gradient operates in the row space with such high magnitude (scaled by $1/\gamma^{2}$) that it dictates the update direction, validating the omission of the prior in the time-marginal step.
6.2 MCMC sampling with warm start
@@
When applying Markov chain Monte Carlo (MCMC), an informative initial state significantly improves both convergence speed and numerical accuracy. As shown in Fig 6, different initialization strategies—including pure-noise initialization, high–order ODE-based initialization, and Tweedie-based estimation—are evaluated under identical MCMC settings (learning rate $1\times 10^{-3}$, two refinement steps, and a $4\times$ downsampling operator). All schemes rapidly approach the measurement-consistent solution; however, the warm start obtained from Tweedie’s estimator produces notably smoother updates with reduced oscillation. After the MCMC refinement step, the Tweedie-initialized trajectories exhibit markedly fewer high-frequency artifacts compared with those initialized via higher–order ODE solvers, demonstrating its advantage as a stable and noise-robust initialization method.
@@
Figure 6: Visualization of 2-step MCMC iterations for $4\times$ downsampling under different initialization setups. Subfigures (a)–(c) are initialized at $\sigma_{\max}=100$ using: (a) Tweedie’s formula, (b) a 5-step Euler ODE solver, and (c) a 5-step RK-4 ODE solver. For comparison, (d) shows standard pure noise initialization at $\sigma=1$.
7 Detailed Algorithm Implementation
Algorithm 1 DAPS++ Inference with E–M Alternating Refinement
0:  Measurement $\mathbf{y}$, operator $\mathcal{A}$, effective noise variance $\gamma^{2}$, diffusion score $s_{\theta}$, noise schedule $\{\sigma_{t}\}$, threshold $\bar{\sigma}$, refinement steps $J$, step sizes $\{\eta_{j}\}$, total E–M cycles $K$.
0:  Final reconstruction $\hat{\mathbf{x}}_{0}$.
1:  Initialize $x_{\mathrm{in}}^{(0)}\sim\mathcal{N}(\mathbf{0},\sigma_{\max}^{2}\mathbf{I})$.
2:  for $k=1$ to $K$ do
3:   /* E-step: One-shot diffusion initialization */
4:   if $\sigma_{\mathrm{in}}>\bar{\sigma}$ then
5:    $\hat{\mathbf{x}}_{0}^{(k)}\leftarrow x_{\mathrm{in}}^{(k-1)}+\sigma_{\mathrm{in}}^{\,2}s_{\theta}\!\big(x_{\mathrm{in}}^{(k-1)},\sigma_{\mathrm{in}}\big)$
6:   else
7:    $\hat{\mathbf{x}}_{0}^{(k)}\leftarrow\texttt{ODESolver}\big(x_{\mathrm{in}}^{(k-1)},\sigma_{\mathrm{in}}\big)$
8:   end if
9:   /* M-step: Likelihood-driven refinement */
10:   $z^{(0)}\leftarrow\hat{\mathbf{x}}_{0}^{(k)}$
11:   for $j=1$ to $J$ do
12:    $\mathbf{r}^{(j-1)}\leftarrow\mathbf{y}-\mathcal{A}(z^{(j-1)})$
13:    $\nabla_{z}\mathcal{L}\leftarrow\frac{1}{\gamma^{2}}\texttt{VJP}_{\mathcal{A}}\big(z^{(j-1)},\mathbf{r}^{(j-1)}\big)$
14:    Sample $\boldsymbol{\xi}_{j}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
15:    $z^{(j)}\leftarrow z^{(j-1)}+\eta_{j}\nabla_{z}\mathcal{L}+\sqrt{2\eta_{j}}\,\boldsymbol{\xi}_{j}$ {ULA refinement}
16:   end for
17:   $\tilde{\mathbf{x}}_{0}^{(k)}\leftarrow z^{(J)}$
18:   /* Re-noising to start next E-step */
19:   Sample $\boldsymbol{\epsilon}_{k}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
20:   $x_{\mathrm{in}}^{(k)}\leftarrow\tilde{\mathbf{x}}_{0}^{(k)}+\sigma_{k}\,\boldsymbol{\epsilon}_{k}$
21:  end for
22:  return $\hat{\mathbf{x}}_{0}\leftarrow\tilde{\mathbf{x}}_{0}^{(K)}$
8 Discussions
8.1 Sampling Efficiency
@@
Figure 7: Phase retrieval results using the top 100 images from the FFHQ test dataset. (a) PSNR, (b) SSIM, and (c) LPIPS statistics (mean $\pm$ standard deviation) show that DAPS++ achieves reconstruction quality comparable to DAPS-4K while requiring substantially fewer sampling steps. (d) Failure rate (PSNR << 25) shows that instability arises from the nonlinear frequency-domain measurement operator.
@@
Sampling efficiency is a critical factor for diffusion-based inverse problem solvers. The computational cost of these methods depends heavily on both the number of neural function evaluations (NFEs) and the refinement steps performed at each iteration. In Tab. 4, we summarize the ODE steps, annealing steps, refinement steps, and resulting NFE budgets for several pixel-space baseline methods and for DAPS under multiple configurations, along with the corresponding single-image sampling times on FFHQ-256, compared against DAPS++. As shown, DAPS++ achieves substantially higher efficiency than existing baselines, delivering competitive performance even with very low NFE budgets.
Configuration ODE/Sampling Steps Annealing Steps Refinement Steps NFE Seconds/Image
@@
Table 4: Sampling time of DAPS on the super-resolution task with FFHQ-256 based on $\bar{\sigma}=0.5$. Reported values denote non-parallel single-image sampling time on the FFHQ-256 dataset using one NVIDIA A100 PCIe 80GB GPU.
8.2 Regarding Phase Retrieval
DAPS has demonstrated strong performance on the phase retrieval task compared with other baselines. However, this advantage is not consistently maintained across the entire dataset. As illustrated in Fig. 7, the variance of failure cases among the top 100 images from the FFHQ test set remains significant. These failures primarily stem from the nonlinear frequency-domain measurement operator, which interacts unfavorably with the pixel-domain initialization produced by the diffusion prior. A trajectory comparison of the diffusion output $\hat{x}_{0}$ in Fig. 8 further shows that DAPS++ exhibits more stable refinement behavior and achieves more reliable reconstructions under the same experimental conditions.
Figure 8: Visualization of the trajectory of diffusion outputs in the phase retrieval task for (a) DAPS-4K and (b) DAPS++-500.
@@ 8.2 Regarding Phase Retrieval
8.3 Future Directions
By decoupling the solution into two iterative and independent stages, our framework avoids the complex requirement of treating the likelihood term $p(\mathbf{y}\mid\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}))$ as an intrinsic part of the diffusion-induced probability manifold. Instead, each stage can be analyzed and optimized in isolation. Specifically, the optimization phase (M-step) is liberated from the constraints of maintaining time-marginal consistency or strictly adhering to the diffusion trajectory. This independence provides a more flexible and theoretically robust foundation for applying diffusion models to measurement-driven reconstruction.
@@ 9 Experimental Details
9.1 Final Refinement via RK4 Solver
@@
$\mathbf{f}(\mathbf{x}_{t},\sigma_{t})=-\dot{\sigma}_{t}\,\sigma_{t}\,s_{\theta}(\mathbf{x}_{t},\sigma_{t}),$ (17)
@@
$\mathbf{x}_{t-h}$ $=\mathbf{x}_{t}+\frac{h}{6}\!\left(k_{1}+2k_{2}+2k_{3}+k_{4}\right).$ (18)
@@
Table 5: Hyperparameter comparison between DAPS and DAPS++. We list the step size $\eta_{0}$, noise scale $\delta$, and total sampling steps. Note that while DAPS requires 100 steps for all tasks, DAPS++ achieves efficient reconstruction with significantly fewer steps (e.g., 2 steps for Super-Resolution).
@@
$t_{i}=t_{\min}+\bigl(t_{1}-t_{\min}\bigr)\left(\frac{i}{N-1}\right)^{\rho}.$ (19)
While standard diffusion training typically utilizes $\rho=7$ and $t_{\min}=0.02$, we tailor the schedule for the inference phase of DAPS++ (specifically with $\gamma=0.05$) to better suit the refinement requirements in the inverse problem space. Accordingly, we set $\rho=-7$ to induce a faster decay of the noise level, thereby concentrating a greater density of discretization steps within the small-noise regime, as illustrated in Fig. 9. With this schedule, given the refinement threshold $\bar{\sigma}=0.5$, approximately 30% of the total sampling steps are allocated to the refinement stage ($\sigma_{t}\leq\bar{\sigma}$), ensuring detailed reconstruction.
Figure 9: Polynomial interpolation schedules. The standard $\rho=7$ (red) maintains high noise levels, whereas our $\rho=-7$ (blue) accelerates decay. This allocates more steps to the small-noise regime, crucial for detailed refinement in inverse problems.
9.2 Hyperparameter Choice
To evaluate the proposed method across general linear and nonlinear inverse problems, we maintain a unified hyperparameter setting to demonstrate robustness. Detailed configurations for all tasks—including Super-Resolution ($4\times$), Inpainting, Deblurring (Gaussian/Motion/Nonlinear), and High Dynamic Range (HDR)—are summarized in Tab. 5, where we provide a direct comparison with the baseline DAPS [zhang2025improving].
Efficiency and NFEs. For the annealing sampling process, we strictly limit computation to a single-step solver per time step. Unlike methods that require multiple gradient steps or extensive MCMC transitions at each noise level, our approach achieves fast and robust reconstruction with minimal computational overhead. As shown in Tab. 5, while the original DAPS requires a fixed budget of $100$ MCMC steps for all tasks, DAPS++ significantly reduces the inference cost without compromising quality.
Annealing Steps for Different Datasets. As reported in Tab. 2 and Tab. 1, we adjust the total number of annealing steps ($K$) based on the dataset complexity. Specifically, we employ a higher number of steps for ImageNet compared to FFHQ. This increase is necessitated by the significantly higher semantic diversity and manifold complexity of ImageNet, which requires a finer discretization of the reverse trajectory to prevent the sampling path from diverging. Conversely, FFHQ, being a domain-specific face dataset with lower entropy, allows for faster convergence with fewer steps.
@@ 9.2 Hyperparameter Choice
$\eta_{t}=\eta_{0}\left[\delta+\frac{t}{T}(1-\delta)\right],$ (20)
@@
10 More Ablation Results
@@
Figure 10: Qualitative results on representative inverse problems on the FFHQ-256 dataset. We compare reconstructions across varying noise refinement threshold $\bar{\sigma}\in\{0.2,0.5,2.0,5.0\}$. The tasks are: (a) Gaussian blur reconstruction; (b) Nonlinear blur reconstruction; (c) High dynamic range (HDR) reconstruction; (d) Motion blur reconstruction; (e) $4\times$ super-resolution; and (f) Box inpainting.
@@
[…]
We present additional ablation studies on the choice of $\bar{\sigma}$ and the polynomial exponent $\rho$ in the noise schedule. Results across several inverse problems are summarized in Tab. 6 and Tab. 7, evaluated under identical measurement noise levels. Illustrative examples are provided in Fig. 10 and Fig. 11.
As shown in Tab. 6, using a larger $\bar{\sigma}$ increases the number of required NFEs but does not yield noticeable improvements in reconstruction quality; in some tasks, it even slightly degrades performance due to excessive early-stage denoising. Conversely, a smaller $\bar{\sigma}$ tends to improve SSIM but may oversuppress fine structures, reducing perceptual fidelity.
For the polynomial schedule, Tab. 7 shows that negative values of $\rho$ consistently enhance performance across tasks, indicating that spending more iterations in the low-noise regime benefits the refinement step. This behavior aligns with our EM-style interpretation: once the diffusion model provides a reliable initialization, allocating more resolution to small-noise updates improves measurement consistency without sacrificing stability.
Figure 11: Qualitative results on representative inverse problems on the FFHQ-256 dataset. We compare reconstructions across varying polynomial interpolation schedules $\rho\in\{-2,-5,-7,2,5,7\}$. The tasks are: (a) Gaussian blur reconstruction; (b) Nonlinear blur reconstruction; (c) High dynamic range (HDR) reconstruction; (d) Motion blur reconstruction; (e) $4\times$ super-resolution; and (f) Box inpainting

### v2

@@ 5 Conclusion
In this work, we revisit the underlying mechanism of diffusion-based inverse problem solvers and propose a fully decoupled initialization-refinement formulation that makes explicit the separation between the diffusion and data-consistency stages. Based on this perspective, we develop DAPS++, a fully decoupled framework that substantially improves sampling efficiency while preserving high reconstruction fidelity. The diffusion stage functions as a prior-driven initialization on a flat manifold, whereas the data-consistency stage performs likelihood-guided refinement independent of the diffusion process. This separation offers a clearer functional interpretation of diffusion-based inference and provides an explanation for the empirical behavior of existing approaches. Across both linear and nonlinear imaging tasks, DAPS++ achieves strong performance using only a fraction of the function evaluations required by conventional methods, establishing an efficient and scalable foundation for future research in diffusion-based inverse problems.
@@
S1.1 Lipschitz Analysis of Prior–Likelihood Interaction
@@
$\mathbf{r}_{t}:=\mathbf{y}-\mathcal{A}\,\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}).$ (S1)
@@
$\kappa_{t}\;\geq\;\frac{\sigma_{\min}^{+}(\mathcal{A})}{\gamma^{2}C}\,\sigma_{t}\,\|\mathbf{r}_{t}\|.$ (S2)
@@
$\|\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t};\sigma_{t})\|\;\lesssim\;\frac{C}{\sigma_{t}}.$ (S3)
@@
$\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}\mid\mathbf{x}_{t};\sigma_{t})=\frac{1}{\gamma^{2}}\mathcal{A}^{\top}\mathbf{r}_{t}+O(\sigma_{t}).$ (S4)
@@
$\|\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}\mid\mathbf{x}_{t};\sigma_{t})\|\;\geq\;\frac{\sigma_{\min}^{+}(\mathcal{A})}{\gamma^{2}}\|\mathbf{r}_{t}\|.$ (S5)
(iii) Lower bound on gradient ratio. Combining (S3) and (S5),
$\kappa_{t}\;\gtrsim\;\frac{\sigma_{\min}^{+}(\mathcal{A})}{\gamma^{2}C}\,\sigma_{t}\,\|\mathbf{r}_{t}\|,$ (S6)
establishing (S2). $\square$
@@
This analysis rigorously justifies the observation that the prior gradient contributes negligibly in high-noise regimes. For standard EDM schedulers in DAPS with $\sigma_{\max}\approx 100$ and low observation noise ($\gamma\approx 0.01$), the gradient ratio reaches magnitudes of $\kappa_{t}\approx 10^{6}$, rendering the prior term mathematically insignificant during early sampling. As illustrated in Fig.˜2, this dominance persists throughout the diffusion process, where the likelihood gradient continues to outweigh the prior contribution even at the minimal noise levels ($\sigma_{\min}\approx 0.1$). Consequently, even for ill-posed problems where $\mathcal{A}$ has a null space, the likelihood gradient operates in the row space with such high magnitude (scaled by $1/\gamma^{2}$) that it dictates the update direction, validating the omission of the prior in the time-marginal step.
S1.2 MCMC Sampling with Warm Start
@@
When applying Markov chain Monte Carlo (MCMC), an informative initial state significantly improves both convergence speed and numerical accuracy. As shown in Fig.˜S1, different initialization strategies (including pure-noise initialization, high-order ODE-based initialization, and Tweedie-based estimation) are evaluated under identical MCMC settings (learning rate $1\times 10^{-3}$, two refinement steps, and a $4\times$ downsampling operator). All schemes rapidly approach the measurement-consistent solution; however, the warm start obtained from Tweedie’s estimator produces notably smoother updates with reduced oscillation. After the MCMC refinement step, the Tweedie-initialized trajectories exhibit markedly fewer high-frequency artifacts compared with those initialized via higher-order ODE solvers, demonstrating its advantage as a stable and noise-robust initialization method.
@@
Figure S1: Visualization of 2-step MCMC iterations for $4\times$ downsampling under different initialization setups. Subfigures (a)–(c) are initialized at $\sigma_{\max}=100$ using: (a) Tweedie’s formula, (b) a 5-step Euler ODE solver, and (c) a 5-step RK-4 ODE solver. For comparison, (d) shows standard pure noise initialization at $\sigma=1$.
S1.3 Convergence Analysis
Building on the assumptions in Sec.˜S1.1, we show that the diffusion initialization provides a warm start of quality $\mathcal{O}(\sigma_{t})$ at every cycle, and that the ULA refinement converges in $\mathcal{O}(\kappa(\mathcal{A})^{2})$ steps.
Initialization quality.
Both methods can be viewed as numerical integrators of the same probability flow ODE from $\sigma_{t}$ to 0. Tweedie’s formula corresponds to a single Euler step with step size $h=\sigma_{t}$ under a EDM diffusion [karras2022elucidating] setting, incurring a truncation error of $\mathcal{O}(\sigma_{t}^{2})$; an $N$-step RK4 solver with step size $h=\sigma_{t}/N$ reduces this to $\mathcal{O}(\sigma_{t}^{4}/N^{4})$. The threshold $\bar{\sigma}$ switches from the single-step Euler (Tweedie) to the higher-order solver when the first-order truncation error begins to exceed the measurement noise floor $\gamma$.
Warm-start convergence of ULA.
Under Assumption A2, the negative log likelihood satisfies $\gamma^{-2}\mathcal{A}^{\top}\mathcal{A}$, which is $\mu$-strongly convex and $L$-smooth when restricted to $\mathrm{Row}(\mathcal{A})$, with
$\mu=\frac{[\sigma_{\min}^{+}(\mathcal{A})]^{2}}{\gamma^{2}},\qquad L=\frac{\|\mathcal{A}\|^{2}}{\gamma^{2}}.$ (S7)
Starting from $\hat{\mathbf{x}}_{0}$ at distance $e_{t_{k}}$ from measurement-consistent solution $\mathbf{x}^{*}$, the step size $\eta\leq 1/L$ satisfies the standard $W_{2}$ convergence bound [10.1111/rssb.12183, AAP1238]:
$W_{2}\!\bigl(\mathrm{Law}(x^{(K)}),\,\pi\bigr)\;\leq\;(1-\eta\mu)^{K}\,e_{t_{k}}\;+\;\mathcal{O}\!\!\left(\sqrt{\frac{\eta\,d_{\mathrm{row}}}{\mu}}\right),$ (S8)
where $\pi\propto\exp(-\ell)$ is the target likelihood distribution and $d_{\mathrm{row}}=\mathrm{rank}(\mathcal{A})$. The first term captures exponential contraction of the initial error; the second is the discretization bias inherent to ULA.
To determine the required step count, we set the contraction term below a target tolerance $\delta$:
$(1-\eta\mu)^{K}\,e_{t_{k}}\;\leq\;\delta\quad\Longrightarrow\quad K\,\ln(1-\eta\mu)\;\leq\;\ln\frac{\delta}{e_{t_{k}}}.$ (S9)
Applying $\ln(1-x)\leq-x$ for $x=\eta\mu\in(0,1)$ and rearranging gives
$K\geq\frac{1}{\eta\mu}\ln\frac{e_{t_{k}}}{\delta}.$ (S10)
Substituting the step size $\eta=1/L$ and noting that $L/\mu=\|\mathcal{A}\|^{2}/[\sigma_{\min}^{+}(\mathcal{A})]^{2}=\kappa(\mathcal{A})^{2}$:
$K\;=\;\mathcal{O}\!\left(\frac{1}{\eta\mu}\,\log\frac{e_{t_{k}}}{\delta}\right)\;=\;\mathcal{O}\!\left(\kappa(\mathcal{A})^{2}\,\log\frac{e_{t_{k}}}{\delta}\right),$ (S11)
In Algorithm 1, re-noising at the end of each cycle adds noise of magnitude $\sigma_{t_{k+1}}$, so corrections below this scale are erased. Setting $\delta=\mathcal{O}(\sigma_{t_{k+1}})$ and substituting $e_{t_{k}}=\mathcal{O}(\sigma_{t_{k}})$:
$K\;=\;\mathcal{O}\!\left(\kappa(\mathcal{A})^{2}\,\log\frac{\sigma_{t_{k}}}{\sigma_{t_{k+1}}}\right)\;=\;\mathcal{O}\!\bigl(\kappa(\mathcal{A})^{2}\,\log\rho\bigr),$ (S12)
where $\rho=\sigma_{t_{k}}/\sigma_{t_{k+1}}$ is the annealing ratio. For geometric schedules, $\rho$ is constant, and the dominant computational factor is $\kappa(\mathcal{A})^{2}$, which is determined by the spectral properties of the forward operator.
Convergence across problem types.
For nonlinear operators, $\kappa(\mathcal{A})$ in (S12) is replaced by the local effective condition number $\kappa_{\mathrm{eff}}=\|\nabla\mathcal{A}(\hat{\mathbf{x}}_{0})\|/\sigma_{\min}^{+}(\nabla\mathcal{A}(\hat{\mathbf{x}}_{0}))$ of the Jacobian at the initialization. This explains the empirical variation in refinement steps: well-conditioned operators such as downsampling ($\kappa_{\mathrm{eff}}$ small, $K=2$) converge rapidly, while operators with broader eigenvalue spread such as Gaussian deblurring require more steps. For strongly nonlinear operators where the local likelihood is non-convex (e.g., phase retrieval), the strong-convexity assumption breaks down and local refinement may converge to an incorrect basin.
S2 Discussions
S2.1 Sampling Efficiency
@@
Figure S2: Phase retrieval results using the top 100 images from the FFHQ test dataset. (a) PSNR, (b) SSIM, and (c) LPIPS statistics (mean $\pm$ standard deviation) show that DAPS++ achieves reconstruction quality comparable to DAPS-4K while requiring substantially fewer sampling steps. (d) Failure rate (PSNR << 25) shows that instability arises from the nonlinear frequency-domain measurement operator.
@@
Sampling efficiency is a critical factor for diffusion-based inverse problem solvers. The computational cost of these methods depends heavily on both the number of neural function evaluations (NFEs) and the refinement steps performed at each iteration. In Tab.˜S1, we summarize the ODE steps, annealing steps, refinement steps, and resulting NFE budgets for several pixel-space baseline methods and for DAPS under multiple configurations, along with the corresponding single-image sampling times on FFHQ-256, compared against DAPS++. As shown, DAPS++ achieves substantially higher efficiency than existing baselines, delivering competitive performance even with very low NFE budgets.
Table S1: Sampling time of DAPS on the super-resolution task with FFHQ-256 based on $\bar{\sigma}=0.5$. Reported values denote non-parallel single-image sampling time on the FFHQ-256 dataset using one NVIDIA A100 PCIe 80GB GPU.
Config. ODE Steps Anneal. Steps Refine. Steps NFE Sec./Img.
@@
S2.2 Regarding Phase Retrieval
DAPS has demonstrated strong performance on the phase retrieval task compared with other baselines. However, this advantage is not consistently maintained across the entire dataset. As illustrated in Fig.˜S2, the variance of failure cases among the top 100 images from the FFHQ test set remains significant. These failures primarily stem from the nonlinear frequency-domain measurement operator, which interacts unfavorably with the pixel-domain initialization produced by the diffusion prior. A trajectory comparison of the diffusion output $\hat{x}_{0}$ in Fig.˜S3 further shows that DAPS++ exhibits more stable refinement behavior and achieves more reliable reconstructions under the same experimental conditions.
Figure S3: Visualization of the trajectory of diffusion outputs in the phase retrieval task for (a) DAPS-4K and (b) DAPS++-500.
@@
S2.3 Future Directions
By decoupling the solution into two iterative and independent stages, our framework avoids the complex requirement of treating the likelihood term $p(\mathbf{y}\mid\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}))$ as an intrinsic part of the diffusion-induced probability manifold. Instead, each stage can be analyzed and optimized in isolation. Specifically, the refinement phase is liberated from the constraints of maintaining time-marginal consistency or strictly adhering to the diffusion trajectory. This independence provides a more flexible and theoretically robust foundation for applying diffusion models to measurement-driven reconstruction.
@@
S3.1 Final Refinement via RK4 Solver
@@
$\mathbf{f}(\mathbf{x}_{t},\sigma_{t})=-\dot{\sigma}_{t}\,\sigma_{t}\,s_{\theta}(\mathbf{x}_{t},\sigma_{t}),$ (S13)
@@
$\mathbf{x}_{t-h}$ $=\mathbf{x}_{t}+\frac{h}{6}\!\left(k_{1}+2k_{2}+2k_{3}+k_{4}\right).$ (S14)
Table S2: Hyperparameter comparison between DAPS and DAPS++. We list the step size $\eta_{0}$, noise scale $\delta$, and total sampling steps. Note that while DAPS requires 100 steps for all tasks, DAPS++ achieves efficient reconstruction with significantly fewer steps (e.g., 2 steps for Super-Resolution).
@@
$t_{i}=t_{\min}+\bigl(t_{1}-t_{\min}\bigr)\left(\frac{i}{N-1}\right)^{\rho}.$ (S15)
While standard diffusion training typically utilizes $\rho=7$ and $t_{\min}=0.02$, we tailor the schedule for the inference phase of DAPS++ (specifically with $\gamma=0.05$) to better suit the refinement requirements in the inverse problem space. Accordingly, we set $\rho=-7$ to induce a faster decay of the noise level, thereby concentrating a greater density of discretization steps within the small-noise regime, as illustrated in Fig.˜S4. With this schedule, given the refinement threshold $\bar{\sigma}=0.5$, approximately 30% of the total sampling steps are allocated to the refinement stage ($\sigma_{t}\leq\bar{\sigma}$), ensuring detailed reconstruction.
Figure S4: Polynomial interpolation schedules. The standard $\rho=7$ (red) maintains high noise levels, whereas our $\rho=-7$ (blue) accelerates decay. This allocates more steps to the small-noise regime, crucial for detailed refinement in inverse problems.
S3.2 Hyperparameter Choice
To evaluate the proposed method across general linear and nonlinear inverse problems, we maintain a unified hyperparameter setting to demonstrate robustness. Detailed configurations for all tasks, which include Super-Resolution ($4\times$), Inpainting, Deblurring (Gaussian/Motion/Nonlinear), and High Dynamic Range (HDR), are summarized in Tab.˜S2, where we provide a direct comparison with the baseline DAPS [zhang2025improving].
Efficiency and NFEs. For the annealing sampling process, we strictly limit computation to a single-step solver per time step. Unlike methods that require multiple gradient steps or extensive MCMC transitions at each noise level, our approach achieves fast and robust reconstruction with minimal computational overhead. As shown in Tab.˜S2, while the original DAPS requires a fixed budget of $100$ MCMC steps for all tasks, DAPS++ significantly reduces the inference cost without compromising quality.
Annealing Steps for Different Datasets. As reported in Tabs.˜2 and 1, we adjust the total number of annealing steps ($K$) based on the dataset complexity. Specifically, we employ a higher number of steps for ImageNet compared to FFHQ. This increase is necessitated by the significantly higher semantic diversity and manifold complexity of ImageNet, which requires a finer discretization of the reverse trajectory to prevent the sampling path from diverging. Conversely, FFHQ, being a domain-specific face dataset with lower entropy, allows for faster convergence with fewer steps.
@@
$\eta_{t}=\eta_{0}\left[\delta+\frac{t}{T}(1-\delta)\right],$ (S16)
@@
S4 More Ablation Results
[…]
@@
Figure S5: Qualitative results on representative inverse problems on the FFHQ-256 dataset. We compare reconstructions across varying noise refinement threshold $\bar{\sigma}\in\{0.2,0.5,2.0,5.0\}$. The tasks are: (a) Gaussian blur reconstruction; (b) Nonlinear blur reconstruction; (c) High dynamic range (HDR) reconstruction; (d) Motion blur reconstruction; (e) $4\times$ super-resolution; and (f) Box inpainting.
@@
We present additional ablation studies on the choice of $\bar{\sigma}$ and the polynomial exponent $\rho$ in the noise schedule. Results across several inverse problems are summarized in Tabs.˜S3 and S4, evaluated under identical measurement noise levels. Illustrative examples are provided in Figs.˜S5 and S6.
As shown in Tab.˜S3, using a larger $\bar{\sigma}$ increases the number of required NFEs but does not yield noticeable improvements in reconstruction quality; in some tasks, it even slightly degrades performance due to excessive early-stage denoising. Conversely, a smaller $\bar{\sigma}$ tends to improve SSIM but may oversuppress fine structures, reducing perceptual fidelity.
For the polynomial schedule, Tab.˜S4 shows that negative values of $\rho$ consistently enhance performance across tasks, indicating that spending more iterations in the low-noise regime benefits the refinement step. This behavior aligns with our decoupled initialization-refinement perspective: once the diffusion model provides a reliable initialization, allocating more resolution to small-noise updates improves measurement consistency without sacrificing stability.
Figure S6: Qualitative results on representative inverse problems on the FFHQ-256 dataset. We compare reconstructions across varying polynomial interpolation schedules $\rho\in\{-2,-5,-7,2,5,7\}$. The tasks are: (a) Gaussian blur reconstruction; (b) Nonlinear blur reconstruction; (c) High dynamic range (HDR) reconstruction; (d) Motion blur reconstruction; (e) $4\times$ super-resolution; and (f) Box inpainting.